Shift report 20030114 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Jan 14 2003 - 00:19:25 EST

  • Next message: Apache: "Shift report 20030114 00:00-08:00"
    ------------------------------------------------------------
    Date: Tue Jan 14 00:19:25 2003, Shift: 16:00-24:00
    Supervisor: cr, with: or
    Log book#13 pages 57 - 59
    
    Shift Summary:
    ------------------------------------------------------------
    1. Machine setup until 21.
    2. 7362 - pedestal run.
       7363 - TPC Pedestals
       7364 - DVM run
       7365 - Data run - finished the 1/4 field A pol (250k trig6)
       7366 - Data run, at 0 field, but unfortunatelly with all magnets.
       7367 - Data run with D1 at 1/4 A pol, to be continued the next shift.
    3. All detectors are working, except for T3.
    
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Tue Jan 14 2003 - 00:25:26 EST